THE A^{2}Π-X^{2}Σ^{+} BAND SYSTEM OF SrF REVISITED

\underline{TIMOTHY C. STEIMLE} , ANH LE, Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287.

Recently DeMille et al^{a} proposed using ^{87}SrF and other heavy diatomic molecules to measure nuclear spin-dependent parity violation (NSD-PV). In this scheme the N=0 (+-parity) levels of the X^{2}Σ^{+} (v=0) state will be magnetically tuned into near degeneracy with the N=1(- parity) levels. The pairs of nearly degenerate levels are mixed by NSD-PV interactions. The process will be monitored using the optical A^{2}Π-X^{2}Σ^{+} transition. Here we report on the analysis of the (1,0) band for ^{88}SrF and the (0,0) band for ^{87}SrF of the A^{2}Π-X^{2}Σ^{+} system recorded at near natural linewidth limit. A combined fit of the optical spectra with the previously recorded rf-transitions^{b} and pure rotational transitions^{c} was performed.

aD. DeMille, S.B. Cahn, D. Murphree, D.A. Rahmlow and M.G. Kozlov, Phys. Rev. Lett. \underline{100} 023003, 2008.

bY. Azuma, W. J. Childs and G. L. Goodman, T. C. Steimle, J. Chem. Phys. \underline{93} 1990 93, 5533.

cH.-U. Schltz-Pahlmann, Ch. Ryzlewicz, J. Hoeft,and T. Trring, 93,74 (1982)Chem. Phys. Lett. \underline{93} 74, 1982.